### v1

# Machine Unlearning Meets Adversarial Robustness via Constrained Interventions on LLMs

Abstract
With the increasing adoption of Large Language Models (LLMs), more customization is needed to ensure privacy-preserving and safe generation. We address this objective from two critical aspects: unlearning of sensitive information and robustness to jail-breaking attacks. We investigate various constrained optimization formulations that address both aspects in a unified manner, by finding the smallest possible interventions on LLM weights that either make a given vocabulary set unreachable or embed the LLM with robustness to tailored attacks by shifting part of the weights to a safer region. Beyond unifying two key properties, this approach contrasts with previous work in that it doesn’t require an oracle classifier that is typically not available or represents a computational overhead. Surprisingly, we find that the simplest point-wise constraint-based intervention we propose leads to better performance than max-min interventions, while having a lower computational cost. Comparison against state-of-the-art defense methods demonstrates superior performance of the proposed approach.
1 Introduction
Ensuring the safety of generative model outputs is a critical requirement for their deployment in real-world and safety-sensitive applications such as online content moderation and confidential data processing. To mitigate the risks of generating toxic or harmful content, prior work has explored strategies such as fine-tuning krause2021gedi ; dengresidual , prompt engineering luo2023prompt ; xie2023defending , and model augmentation dai2022knowledge ; meng2022locating . However, fine-tuning and model augmentation are computationally expensive, while prompt-based defenses are often brittle and sensitive to changes in model, context, and adversarial manipulation.
A more efficient alternative leverages lightweight probes (e.g., classifiers) on the activation space to steer generations toward desired attributes such as topic or sentiment dathathriplug ; subramani2022extracting ; hernandezinspecting ; konen2024style ; cheng2024linearly . While effective at reducing computational overhead, these methods remain heavily constrained by the probe’s limited training data and are ineffective in defending against adversarial attacks. Recent works have instead turned to adversarial detection, analyzing statistical signatures of perturbed inputs or comparing them against safe baselines candogansingle2025 ; robeysmoothllm2025 .
Conversely, the task of unlearning sensitive or confidential content has mainly been addressed through partial retraining in teacher–student frameworks chen2023unlearn or iterative fine-tuning yao2024large ; tian2024forget , both of which incur substantial computational cost. Thus, there remains a gap for methods that jointly address robustness and unlearning while remaining computationally efficient. In this work, we bridge this gap by introducing a simple yet effective approach that simultaneously tackles adversarial robustness and unlearning. Specifically, our contributions can be summarized as follows:
  • We propose and solve various constrained optimization problems to embed LLMs with robustness to adversarial attacks, while allowing them to unlearn unwanted content,
  • We overcome the need for artificial probes by introducing a continuous relaxation over the space of prompts and performing interventions that directly constrain concept embeddings,
  • We demonstrate a gain in performance in comparison to state-of-the-art defense algorithms and set a state-of-the-art for affordable unlearning on LLMs.
2 Problem Setting
Given a trained language model $\ell:\Sigma\longrightarrow\Sigma$, we consider two fundamental safety-related tasks:
  1. 1.
    How to delete certain information, instantiated as a given set of words, from the generation space of $\ell$ at a minimal computational cost ?
  2. 2.
    How to make $\ell$ more robust to jail-breaking adversarial attacks that lead to dangerous or toxic content ?
To address both questions of unlearning and robustness in a unified way, we leverage the fact that information is implicitly stored in latent space pathways, hence constraining them by partially steering weights simultaneously removes both sensitive and toxic content. To achieve that, several strategies warrant attention depending on whether one chooses to maximize distance from risky regions, minimize distance from safe ones or target specific concepts in a point-wise manner. We present in the next section the formal constrained optimization problem descriptions for each of these strategies and discuss how to solve them.
3 Constrained Model Interventions
We present in the following three constrained intervention approaches we propose, as well as the related strengths and drawbacks of each approach.
3.1 Towards Safer Regions (TSR)
In the first formulation, we seek to find minimal weight perturbations that maximize the probability of safe responses to jail-breaking prompts
$\min_{\|{\delta}\|\leq\varepsilon}\quad\mathcal{L}_{\textrm{safety}}(\ell_{\boldsymbol{\theta}+\boldsymbol{\delta}}(\mathbf{x}),\mathbf{y}_{\textrm{safe}}),\\$ (1)
where $\mathcal{L}_{\textrm{safety}}$ represents the safety-oriented loss function that encourages the generation of appropriate refusal responses, $\theta\in\mathbb{R}^{d}$ the model parameters, $\mathbf{y}_{\textrm{safe}}$ the target safe outputs, $x$ a given jailbreaking input embedding and $\varepsilon>0$ the perturbation budget that constrains the magnitude of weight modifications.
The safety loss function is defined as:
$\mathcal{L}_{\textrm{safety}}(f_{\boldsymbol{\theta}+\boldsymbol{\delta}}(\mathbf{x}),\mathbf{y}_{\textrm{safe}})=-\log\left(\sum_{k\in\mathcal{K}_{\textrm{safety}}}p_{k}(\mathbf{x};\boldsymbol{\theta}+\boldsymbol{\delta})\right),$ (2)
where $\mathcal{K}_{\textrm{safety}}$ is the set of token indices corresponding to safety keywords (e.g., "I cannot", "I apologize", "inappropriate"…), and $p_{k}(\mathbf{x};\boldsymbol{\theta}+\boldsymbol{\delta})$ represents the probability assigned to token $k$ by the perturbed model, for input prompt $x$. Among the advantages of this formulation is that, one does not need examples of dangerous generation, which requires annotation. Furthermore, it can simply be solved via projected-gradient descent. However, the model output embeddings are constrained towards safe generation in a soft way, which leads to weaker performance compared to the hard constraint formulation presented in subsection 3.3 below.
3.2 Away from Risky Regions (ARR)
Given the popularity min-max training in adversarial defense on images madry2018towards ; robeyadversarial2024 , we explore such a formulation in the context of LLMs. Specifically, we consider the max-min problem
$\max_{\|\boldsymbol{\delta}\|\leq\varepsilon}\,\,\min_{\mathbf{x}\in\mathcal{X}}$ $\mathcal{L}_{\text{harmful}}(\ell_{\boldsymbol{\theta}+\boldsymbol{\delta}}(\mathbf{x}),\mathbf{y}_{\text{harmful}}).$ (3)
This formulation seeks weight perturbations that minimize the likelihood of generating harmful content, under worst-case input scenarios $x\in\mathcal{X}$, where $\mathcal{X}$ is the prompt space. The harmful loss function is defined as:
$\mathcal{L}_{\text{harmful}}(\ell_{\boldsymbol{\theta}+\boldsymbol{\delta}}(\mathbf{x}),\mathbf{y}_{\text{harmful}})=-\log\left(\sum_{k\in\mathcal{K}_{\text{harmful}}}p_{k}(\mathbf{x};\boldsymbol{\theta}+\boldsymbol{\delta})\right),$ (4)
where $\mathcal{K}_{\text{harmful}}$ contains token indices associated with harmful instruction-following behavior, and $x\in\mathcal{X}$ is an input. Given the discrete structure of $\mathcal{X}$, one needs a continuous relaxation to solve problem (3) efficiently. For that matter, use a probabilistic relaxation over the number of tokens, leading to an optimization over the simplex which can be solved via projected gradient descent geisler2024attacking . Although this approach might be more principled, it requires knowledge or choice of a set of harmful concepts and might be more conservative, given that it considers worst-case prompts.
3.3 Point-Wise Constrained Regions (PCR)
Last, we explore a simple point-wise constraint strategy that is based on finding the smallest intervention that makes the LLM MLP activations, for a jailbreaking prompt, non-equal to embedding of dangerous output. More precisely, denoting by $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ a set of forbidden concept embeddings and by $o^{(l)}(x,\theta)$ the output of layer $l\leq L$ corresponding to input prompt $x$ and parameters $\theta$, the goal is to solve for each $l\leq L$,
[…]
where we abuse notation to mean update in the parameters corresponding to the $l$-th layer of the MLP when writing $\theta+\delta^{l}$. The solution can then be written under closed form based on the Karush–Kuhn–Tucker (KKT) conditions. We report it in Appendix C. Similar to the previous strategy, the main drawback of this formulation is the fact that it requires a set of forbidden concepts $\mathbf{C}$. Yet, given its simplicity, it allows for a semi-closed-form solution which, surprisingly, leads to the best performance even exceeding state-of-the-art defense approaches, as illustrated in section 4.
4 Experimental Results
We evaluate the various intervention strategies on a custom obedience dataset (reported in Appendix B) proposed by geisler2024attacking as well as on the standard LLM defense benchmark HarmBench mazeika2024harmbench . For robustness evaluation, we compare against state-of-the-art adversarial defense algorithms SmoothLLM robeysmoothllm2025 and Self-reminder xie2023defending . We measure performance via adversarial attack success rate (ASR) as well as refusal levels (see appendix A for details). We compute the attacks using the state-of-the-art LLM jailbreaking algorithm proposed by geisler2024attacking . We report running times in Appendix A. As for unlearning levels, we measure them by comparing perplexity levels of given sequences before and post-intervention. Since perplexity is a proxy for how likely a sequence is, this represents a practical measure for unlearning levels. Due to the lack of affordable unlearning algorithms for LLMs, we simply compare the models against themselves post-intervention.
4.1 Adversarial Robustness
The results shown in Tables 1 and 2 demonstrate that our point-wise intervention approach has higher performance than the competitive approaches. As for Table 3, it shows that formulations 1 and 2 barely improve over the unprotected baseline.
Table 1: Attack Success Rate with different defenses - HarmBench Dataset- Lower better
Model No Defense PointWise (Ours) SmoothLLM Self-Reminder
Basic Warn Praise
Llama-3.1 8B 11.0 0.0 7.245 2.03 1.98 0.8
Mistral 7B v0.2 30.0 5.88 18.9 23.1 19.8 28.5
Gemma 2B-IT 22.0 2.508 8.225 12.98 12.76 19.58
Table 2: Refusal Pattern Analysis in Protected Completions - Higher better
Model Our intervention (%) SmoothLLM (%) Self Reminder (%)
Llama-3.1 8B Instruct 100.0 87.5 24.3
Gemma 2B-IT 97.4 10 36.9
Mistral 7B v0.2 26.7 37.5 20
Table 3: Attack success rate across models and formulations (One layer)
Model No Defense PointWise (Ours) SmoothLLM Self-Reminder
Basic Warn Praise
Llama-3.1 8B 11.0 0.0 7.245 2.03 1.98 0.8
Mistral 7B v0.2 30.0 5.88 18.9 23.1 19.8 28.5
Gemma 2B-IT 22.0 2.508 8.225 12.98 12.76 19.58
4.2 Machine Unlearning
Once again, the numerical results reported in Table 4 show the effectiveness of the point-wise intervention strategy in unlearning given sets of words or concepts at a much reduced computational cost.
Table 4: Forbidden Word Perplexity Analysis: Baseline vs. Intervention Results
Model Dataset Baseline PointWise Intervention
Gemma-2B-IT Obedience 8.816 12.72
Gemma-2B-IT HarmBench 16.757 18.157
Mistral-7B-Instruct Obedience 8.627 13.74
Mistral-7B-Instruct HarmBench 6.195 8.005
Llama-3-8B-Instruct Obedience 6.48 7.735
Llama-3-8B-Instruct HarmBench 10.08 13
5 Conclusion
In this work, we introduced a constrained optimization framework that jointly addresses the challenges of unlearning and robustness in large language models, enabling more reliable and effective model customization. By formulating these objectives leveraging relaxations over the space of prompts and tailored point-wise constraints, our approach reduces reliance on external probes, which are often brittle and require substantial computational resources to train at scale. This framework provides a lightweight yet principled alternative, paving the way for safer and more efficient deployment of generative models in real-world settings, especially for smaller organizations and communities with limited resources.
Broader Impact Statement
This work advances the safety of generative models by introducing a computationally efficient framework for robustness and unlearning. In principle, such methods can reduce the risk of harmful outputs (e.g., toxic, biased, or confidential content) and thereby enable safer integration of generative AI into sensitive domains such as education, healthcare, and online platforms. At the same time, more powerful unlearning techniques could be misapplied—for instance, to obscure accountability or selectively erase undesirable information. Careful governance and responsible deployment will therefore be essential to ensure that these methods are used to promote safety and accessibility without introducing new harms.
Acknowlegments
This work was supported by the Swiss National Science Foundation under grant No. 212876. We acknowledge computational resources from the Swiss National Supercomputing Centre CSCS. The authors acknowledge Prof. Hossein Gorji for valuable discussions and insightful feedback. R.D. acknowledges Dr. Ivan Lunati for providing laboratory infrastructure and computational resources.
Appendix A Additional Numerical Results & Experimental Setup
Performance Measures. Attack success rate only captures whether the exact word or sentence have been generated. As for refusal levels, they measure when the model completely refuses to respond, making it a stronger measure of performance.
Performance Layer Scaling. We report in Figures 1 and 2 the evolution of performance as more MLP layers used in interventions.
HarmBench Dataset - Intervention Scope Analysis:
Figure 1: Intervention Configuration Comparison - HarmBench Dataset
Figure 2: Intervention Configuration Comparison - Custom Obedience Dataset
Running Times.
Table 5: Defense Method Timing Comparison (Average per test case)
Model Attack Time (s) Our PCR approach (s) SmoothLLM(s)
LLaMA 3 8B 38.89 20.16 36.12
Mistral-7B-Instruct-v0.2 27.43 17.28 40.17
Gemma 2B-it 14.375 10.44 11.62
Note: Self-Reminder execution times were not recorded as this method introduces negligible computational overhead. The approach operates solely through augmentation of the system prompt with reminder text, requiring no additional forward passes or model modifications during inference.
Appendix B Obedience Dataset
We generate the following forbidden keywords as a complementary dataset: "abuse", "accused", "aggressive", "annihilate", "artillery", "assassinate", "assault", "attack", "battle", "blade", "bleed", "blood", "bomb", "brutality", "cannon", "carnage", "chaos", "combat", "conflict", "crime", "criminal", "cruelty", "culprit", "damage", "deadly", "death", "defendant", "demolish", "destroy", "detonation", "discipline", "disorder", "dynamite", "eliminate", "execute", "explosion", "explosive", "fatal", "fear", "felon", "fight", "frighten", "gore", "grenade", "guilty", "gun", "harm", "horror", "hurt", "injure", "intimidate", "invasion", "judgment", "justice", "kill", "knife", "massacre", "mayhem", "menace", "merciless", "missile", "murder", "nightmare", "obliterate", "offender", "panic", "penalty", "perpetrator", "pistol", "poison", "punishment", "raid", "rebellion", "retaliation", "revenge", "revolt", "rifle", "rocket", "ruthless", "savage", "scare", "sentence", "shoot", "slash", "slaughter", "stab", "suspect", "sword", "terror", "terrorize", "threat", "torture", "toxic", "vengeance", "vicious", "violence", "violent", "war", "warfare", "wound".
Appendix C Closed-form Solution for Point-Wise Interventions
We write down the KKT conditions for the third formulation of the robustness problem. Recall
$\min_{\delta\in\mathbb{R}^{d_{l}}}$ $\|\delta\|_{2}^{2}$
s.t. $\|\,o^{(l)}(x;\theta+\delta)-c_{i}\,\|_{2}\;\;\geq\;\varepsilon,\quad i=1,\dots,n.$
The Lagrangian is given by:
$\mathcal{L}(\delta,\lambda)=\|\delta\|_{2}^{2}+\sum_{i=1}^{n}\lambda_{i}\Big(\varepsilon-\|\,o^{(l)}(x;\theta+\delta)-c_{i}\,\|_{2}\Big),\quad\lambda_{i}\geq 0.$
Leading to KKT Conditions
(Primal feasibility) $\|\,o^{(l)}(x;\theta+\delta^{\star})-c_{i}\,\|_{2}\;\;\geq\;\varepsilon,$ $i=1,\dots,n,$
(Dual feasibility) $\lambda_{i}^{\star}\;\;\geq 0,$ $i=1,\dots,n,$
(Complementary slackness) $\lambda_{i}^{\star}\Big(\varepsilon-\|\,o^{(l)}(x;\theta+\delta^{\star})-c_{i}\,\|_{2}\Big)=0,$ $i=1,\dots,n,$
(Stationarity) $\nabla_{\delta}\mathcal{L}(\delta^{\star},\lambda^{\star})=2\delta^{\star}-\sum_{i=1}^{n}\lambda_{i}^{\star}\,J_{i}(\delta^{\star})^{\top}\frac{o^{(l)}(x;\theta+\delta^{\star})-c_{i}}{\|\,o^{(l)}(x;\theta+\delta^{\star})-c_{i}\,\|_{2}}=0,$
where
$J_{i}(\delta):=\frac{\partial o^{(l)}(x;\theta+\delta)}{\partial\theta^{(l)}}$
Single-Constraint Building Block: For a single forbidden concept, the closed-form solution serves as the fundamental building block:
$\delta^{(l)*}_{\text{single}}=\frac{[\epsilon-\|r_{i}\|_{2}]_{+}}{\|h_{\text{intermediate}}\|_{2}^{2}}\frac{r_{i}h_{\text{intermediate}}^{T}}{\|r_{i}\|_{2}}$ (6)
where $\cdot]_{+}=\max(0,\cdot)$ denotes the positive part function and $h_{\text{intermediate}}$ is the output of previous layer l1l-1.
Most Violated Constraint Approach: For the multi-constraint case, we employ a practical approach based on constraint violation priority. The violation severity for each concept is measured as:
$\text{violation}_{i}=\max(0,\epsilon-\|r_{i}\|_{2})$ (7)
The most violated constraint is identified as:
$j^{*}=\arg\max_{i}\text{violation}_{i}=\arg\min_{i}\|r_{i}\|_{2}\quad\text{(among violated constraints)}$ (8)
Iterative Multi-Constraint Solution: The complete solution employs an iterative refinement algorithm:
Algorithm: Multi-Constraint KKT Solution (9)
$\text{1. Initialize: }\Delta\theta^{(l)}=\mathbf{0}$
$\text{2. For }k=1,2,\ldots,K_{\max}:$
$\quad\text{a. Update current state: }\theta_{\text{current}}^{(l)}=\theta^{(l)}+\Delta\theta^{(l)}$
$\quad\text{b. Compute residuals: }r_{i}^{(k)}=\theta_{\text{current}}^{(l)}h_{\text{intermediate}}-c_{i}$
$\quad\text{c. Find most violated: }j^{*}=\arg\min_{i}\|r_{i}^{(k)}\|_{2}\text{ s.t. }\|r_{i}^{(k)}\|_{2}<\epsilon$
d. If no violations, terminate
$\quad\text{e. Compute single-constraint solution for }c_{j^{*}}$
$\quad\text{f. Apply damped update: }\Delta\theta^{(l)}\leftarrow\Delta\theta^{(l)}+\alpha\Delta\theta_{j^{*}}^{(l)}$
$\text{3. Return }\Delta\theta^{(l)*}$

### v2

@@
 Ensuring the safety of generative model outputs is a critical requirement for their deployment in real-world and safety-sensitive applications such as online content moderation and confidential data processing.
-To mitigate the risks of generating toxic or harmful content, prior work has explored strategies such as fine-tuning krause2021gedi ; dengresidual , prompt engineering luo2023prompt ; xie2023defending , and model augmentation dai2022knowledge ; meng2022locating .
+To mitigate the risks of generating toxic or harmful content, prior work has explored strategies such as fine-tuning krause2021gedi ; dengresidual , prompt engineering luo2023prompt ; xie2023defending , uncertainty quantification dakhmouche2025can and model augmentation dai2022knowledge ; meng2022locating .
 However, fine-tuning and model augmentation are computationally expensive, while prompt-based defenses are often brittle and sensitive to changes in model, context, and adversarial manipulation.
@@
 Thus, there remains a gap for methods that jointly address robustness and unlearning while remaining computationally efficient.
-In this work, we bridge this gap by introducing a simple yet effective approach that simultaneously tackles adversarial robustness and unlearning.
+In this work, inspired by principled robustness approaches for regression dakhmouche2025noise; dakhmouche2025network; dakhmoucherobust, we bridge this gap by introducing a simple yet effective approach that simultaneously tackles adversarial robustness and unlearning.
 Specifically, our contributions can be summarized as follows:
@@
 Mistral 7B v0.2 26.7 37.5 20
-Table 3: Attack success rate across models and formulations (One layer)
+Table 3: Attack Success Rate with different defenses - Custom Dataset- Lower better
 Model No Defense PointWise (Ours) SmoothLLM Self-Reminder
 Basic Warn Praise
-Llama-3.1 8B 11.0 0.0 7.245 2.03 1.98 0.8
-Mistral 7B v0.2 30.0 5.88 18.9 23.1 19.8 28.5
-Gemma 2B-IT 22.0 2.508 8.225 12.98 12.76 19.58
+Llama-3.1 8B 60.50 5.49 30.17 3.00 2.00 18.00
+Mistral 7B v0.2 90.00 36.47 61.00 77.00 66.00 95.00
+Gemma 2B-IT 85.30 30.50 56.20 59.00 58.00 89.00
 4.2 Machine Unlearning
@@
 Careful governance and responsible deployment will therefore be essential to ensure that these methods are used to promote safety and accessibility without introducing new harms.
-Acknowlegments
+Acknowledgements
 This work was supported by the Swiss National Science Foundation under grant No.